# Enhancing Training Efficiency Using Packing with Flash Attention

### v1

@@ 3.5 Example
The modifications required are lightweight and are limited to providing the position IDs to Flash Attention. For details, see the repository at [2]. This solution relies, however, on the model exposing position IDs. As of the time of writing, 13 models expose position IDs and are supported by the solution. Specifically: dbrx, falcon, gemma, llama, mistral, olmo, phi, phi3, qwen2, qwen2_moe, stablelm, starcoder2 are all supported by the solution.
@@ 4 Experiments and Results
We consider three different datasets with different example-length characteristics: FLAN [15], OrcaMath[13], and python code data from The Stack[11]. A subset of 20K examples from each of the 3 datasets was used in the experiments. We evaluate the throughput in tokens/second, denoted Tok/s, peak memory, denoted Mem, and validation loss, denoted VLoss. All tests were performed on a single A100-80GB node having 8 GPU with FSDP.
@@ 4.1 Main Results
We demonstrate the benefits of the solution on 9 of the 13 supported models. Specifically, we test it on Llama-2-7B-fp16, mistralai/Mistral-7B-v0.1, tiiuae/falcon-7b, google/gemma-7b, microsoft/phi-2, Qwen/CodeQwen1.5-7B, bigcode/starcoder2-7b, stabilityai/stablelm-2-1_6b, Qwen/Qwen1.5-MoE-A2.7B.
@@
Observe that our Packing with PositionID methods improve substantially the throughput in terms of tokens per second, in many case above and beyond even basic packing that does not account for improper cross-attention (the second row in each model section).
However, this maximal packing has an impact on the loss behaviour. Due to the fact that far fewer optimisation steps are taken with such maximal packing, the loss does not decrease as fast, and its effect is confirmed by the validation loss ("VLoss") after one epoch. Hence, as a best-of-both worlds remedy, we propose the online mini-batch approach to Packing with PositionIDs, which due to the packing of fewer examples in each pack, does not achieve the maximal throughput, however it achieves the same optimal loss pattern and hence validation loss, as the inefficient padding-based approach. It also improves on throughput substantially beyond the padding-based approach.
The variation in benefit from our proposed approach across the FLAN and OrcaMath datasets can be attributed to the statistical dataset characteristics of each; the very small sample lengths of FLAN, whose primary mode is around 100 tokens, is highly amenable to significant throughput increase from packing, more so than OrcaMath, whose mode is around 400 tokens, as seen in Figure 1. The benefits across model architectures are consistent, with the exception of Gemma-7B and Qwen1.5-MoE-A2.7B.
@@
Table 2: Summary of main results on a 20K sample of the FLAN dataset over 9 of the 13 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online mini-batch.
@@
Table 3: Summary of main results on a 20K sample of the OrcaMath dataset over 9 of the 13 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online mini-batch.
4.2 Ablation Study on Mini-batch Size
In this ablation study we compare the effect of mini-batch size (bs) on training throughput and memory for two approaches: the standard padding vs proposed mini-batch packing. For this experiment, we fine-tune Mistral-7B[10] on 2 different training datasets: FLAN_20k, (right) OrcaMath_20k. We present the throughput results in Figure 3 and the memory results in 2. The advantage of the mini-batch packing is significant, not only is the throughput a factor of 2x faster and the peak memory usage significantly lower than using padding, but the training runs out of memory (OOM) much later allowing for larger batch sizes.
Figure 2: Mini-batch packing offers lower GPU peak memory usage than padding approach as mini-batch size increases. “OOM" denotes out of memory error. Lower is better.
Figure 3: Mini-batch packing offers better throughput (Tokens/s) than padding as mini-batch size increases. “OOM" denotes out of memory error. Higher is better.
@@ 4.3 Ablation Study on Sample Selection Method
In this ablation study we consider two different models: Mistral-7B[10] and CodeLlama-Python-7B [14]. In Tables 4,5, and 6, we examine different batch sizes and sequence lengths on two of the most popular models, Mistral-7B and CodeLlama-Python-7B with several sample selection methods. We provide the results on FLAN, OrcaMath, and the Stack [11], respectively. In the tables, bs is the mini-batch size per GPU, msl stands for maximum sequence length of an example/packed sequence, and gas denotes the number gradient accumulation steps.
As expected, the padding-based approaches have the lowest training throughput. The benefits of any of the packing methods are greater when the dataset sample lengths are smaller, so that, as seen in Tables 4 and 5, FLAN offers the most potential for throughput improvement, then OrcaMath. Here we also examine packing on the python code dataset the Stack, which as we have seen in Figure 1 is characterised by very long sequence lengths in the dataset, and hence offers less througput benefit with packing. Table 6 shows that while our packing with PositionIDs still improves the throughput, the benefits are as expected less substantial than one sees on the datasets with smaller sequences: FLAN, and OrcaMath.
It is instructive to delve into both training and validation loss. We observe that the proposed packing approaches all offer on par or better loss behaviour than the baselines.
@@
Table 4: Finetuning Mistral-7B for 1 epoch on FLAN_20k dataset: performance comparison of different sample packing approaches. Throughput in train tokens/second is improved substantially using the proposed Packing with Position IDs without loss performance degradation.
@@
Table 5: Finetuning Mistral-7B for 1 epoch on OrcaMath_20k dataset: performance comparison of different sample packing approaches. Throughput in train tokens/second is improved substantially using the proposed Packing with Position IDs without loss performance degradation.
@@
Table 6: Finetuning CodeLlama-Python-7B for 1 epoch on the Stack_20k dataset: performance comparison of different sample packing approaches. Throughput in train tokens/second is improved substantially using the proposed Packing with Position IDs without loss performance degradation.

### v3

@@ 3.5 Example
The modifications required are lightweight and are limited to providing the position IDs to Flash Attention. For details, see the repository at [2]. This solution relies, however, on the model exposing position IDs. As of the time of writing, 14 models expose position IDs and are supported by the solution. Specifically: llama, mistral, mixtral, granite, dbrx, falcon, gemma, olmo, phi/phi2, phi3, qwen2, qwen2_moe, stablelm, starcoder2 are all supported by the solution.
@@ 4 Experiments and Results
We consider three different datasets with different example-length characteristics: FLAN [15], OrcaMath[13], and python code data from The Stack[11]. A subset of 20K examples from each of the 3 datasets was used in the experiments. We evaluate the throughput in tokens/second, denoted Tok/s, peak memory, denoted Mem, and validation loss, denoted VLoss. All tests were performed on a single A100-80GB node having 8 GPU with FSDP. The throughput measurements are averaged over the 8 GPU and the peak memory recorded is the maximum over the 8 GPU.
@@ 4.1 Main Results
We demonstrate the benefits of the solution on 10 of the 14 supported models. Specifically, we test it on Llama-2-7B-fp16, mistralai/Mistral-7B-v0.1, granite-8b-code-base, tiiuae/falcon-7b, google/gemma-7b, microsoft/phi-2, Qwen/CodeQwen1.5-7B, bigcode/starcoder2-7b, stabilityai/stablelm-2-1_6b, Qwen/Qwen1.5-MoE-A2.7B.
@@
Observe that the Packing with PositionID methods improve substantially the throughput in terms of tokens per second, in many case above and beyond even basic packing that does not account for improper cross-attention (the second row in each model section).
However, this maximal packing has an impact on the loss behaviour. Due to the fact that far fewer optimisation steps are taken with such maximal packing, the loss does not decrease as fast, and its effect is confirmed by the validation loss ("VLoss") after one epoch. Hence, as a best-of-both worlds remedy, we propose the online minibatch approach to Packing with PositionIDs, which due to the packing of fewer examples in each pack, does not achieve the maximal throughput, however it achieves the same optimal loss pattern and hence validation loss, as the inefficient padding-based approach. It also improves on throughput substantially beyond the padding-based approach.
The variation in benefit from our proposed approach across the FLAN and OrcaMath datasets can be attributed to the statistical dataset characteristics of each; the very small sample lengths of FLAN, whose primary mode is around 100 tokens, is highly amenable to significant throughput increase from packing, more so than OrcaMath, whose mode is around 400 tokens, as seen in Figure 1. This means that offline packing can achieve very significant throughput improvements. On the other hand, with minibatch packing, the padding baseline pads only to the length of the longest sequence in each minibatch. Hence, the variance (rather than the mean) of the sample lengths affects the gain in throughput, with high variance leading to more minibatch padding. This is also confirmed by the higher improvement in throughput from minibatch packing with PositionIDs compared to padding in FLAN, Table 2 versus OrcaMath, in Table 3. The benefits across model architectures are consistent, with the exception of Gemma-7B and Qwen1.5-MoE-A2.7B.
@@
Table 2: Summary of main results on a 20K sample of the FLAN dataset over 10 of the 14 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online mini-batch.
@@
Table 3: Summary of main results on a 20K sample of the OrcaMath dataset over 10 of the 14 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online minibatch.
4.2 Ablation Study on Minibatch Size
In this ablation study we compare the effect of minibatch size (bs) on training throughput and memory for two approaches: the standard padding vs proposed minibatch packing. For this experiment, we fine-tune Mistral-7B[10] on 2 different training datasets: FLAN_20k, (right) OrcaMath_20k. We present the throughput results in Figure 3 and the memory results in 2. The advantage of minibatch packing is significant, not only is the throughput a factor of 2x faster and the peak memory usage significantly lower than using padding, but the training runs out of memory (OOM) much later allowing for larger batch sizes.
Figure 2: Minibatch packing offers lower GPU peak memory usage than padding approach as mini-batch size increases. “OOM" denotes out of memory error. Lower is better.
Figure 3: Minibatch packing offers better throughput (Tokens/s) than padding as mini-batch size increases. “OOM" denotes out of memory error. Higher is better.
@@ 4.3 Ablation Study on Sample Selection Method
In this ablation study we consider two different models: Mistral-7B[10] and CodeLlama-Python-7B [14]. In Tables 4,5, and 6, we examine different batch sizes and sequence lengths on two of the most popular models, Mistral-7B and CodeLlama-Python-7B with several sample selection methods. We provide the results on FLAN, OrcaMath, and the Stack [11], respectively. In the tables, bs is the minibatch size per GPU, msl stands for maximum sequence length of an example/packed sequence, and gas denotes the number gradient accumulation steps. Note that the Stack dataset has high variance in sequence length as seen in Figure 1, and long sequence lengths.
As expected, the padding-based approaches have the lowest training throughput, and offline packing offers the highest throughput. Also as expected and as seen previously minibatch packing offers an intermediate ground with still-signicant throughput increases.
It is instructive to delve into both training and validation loss. The offline packing methods, even using position IDs fail to provide the same level of validation loss after one epoch that padding and minibatch packing with position IDs provide. As noted above, this is due primarily to the fewer number of optimisation update steps performed.
@@
Table 4: Finetuning Mistral-7B for 1 epoch on FLAN_20k dataset: performance comparison of different sample packing approaches. The first group in each section uses minibatches and the second and third groups use offline packing. Throughput (train tokens/second) is improved substantially using Packing with Position IDs. Loss performance is the same for padding and minibatch packing.
@@
Table 5: Finetuning Mistral-7B for 1 epoch on OrcaMath_20k dataset: performance comparison of different sample packing approaches. The first group in each section uses minibatches and the second and third groups use offline packing. Throughput (train tokens/second) is improved substantially using Packing with Position IDs. Loss performance is the same for padding and minibatch packing.
@@
Table 6: Finetuning CodeLlama-Python-7B for 1 epoch on the Stack_20k dataset: performance comparison of different sample packing approaches. The first group in each section uses minibatches and the second and third groups use offline packing. Throughput (train tokens/second) is improved substantially using Packing with Position IDs. Loss performance is the same for padding and minibatch packing.